### v1

# Revisiting Direct Speech-to-Text Translation with Speech LLMs: Better Scaling than CoT Prompting?

Abstract
Recent work on Speech-to-Text Translation (S2TT) has focused on LLM-based models, introducing the increasingly adopted Chain-of-Thought (CoT) prompting, where the model is guided to first transcribe the speech and then translate it. CoT typically outperforms direct prompting primarily because it can exploit abundant Automatic Speech Recognition (ASR) and Text-to-Text Translation (T2TT) datasets to explicitly model its steps. In this paper, we systematically compare CoT and Direct prompting under increasing amounts of S2TT data. To this end, we pseudo-label an ASR corpus by translating its transcriptions into six European languages, and train LLM-based S2TT systems with both prompting strategies at different data scales. Our results show that Direct improves more consistently as the amount of data increases, suggesting that it may become a more effective approach as larger S2TT resources are created.
Index Terms—  speech-to-text translation, speech large language models
1 Introduction
End-to-End (E2E) S2TT is increasingly preferred over the traditional cascaded pipeline of ASR followed by T2TT. Unlike cascaded approaches, E2E systems avoid error propagation across modules and can in principle exploit acoustic and prosodic cues that are lost in intermediate transcripts. Recent studies show that the performance of E2E systems is approaching that of cascaded methods, with some works even reporting improvements beyond them [1].
A popular strategy for building E2E systems is to adapt a backbone LLM already fine-tuned for T2TT to the speech modality [2, 3]. Within this framework, recent works adopt CoT prompting, where the model is guided to first transcribe and then translate [4, 5, 6]. CoT has gained attention because it outperforms Direct translation (without intermediate steps). More importantly, CoT allows to leverage widely available ASR and T2TT data by explicitly modeling these steps.
In this work, we revisit this comparison and argue that the advantage of CoT is mainly tied to the large quantity of ASR and T2TT data currently used for training. We show that as S2TT data becomes increasingly available, direct translation may emerge as a stronger alternative. To the best of our knowledge, this is the only systematic study that compares Direct and CoT prompting under large-scale pseudo-labeled S2TT data with this much language coverage.
The dominant trend for creating S2TT resources is to translate existing ASR corpora into target languages [7, 8]. Following this approach, we construct a pseudo-labeled dataset through a translation pipeline (Figure 1), and train models with both Direct and CoT while scaling the amount of data. Our results demonstrate that Direct exhibits more stable scaling than CoT.
Fig. 1: Overview of the pseudo-labeling pipeline.
This finding mirrors the evolution of ASR itself, where early phoneme- or grapheme-based pipelines were replaced by E2E models that directly map speech to text [9]. We presume that, by enforcing an intermediate transcription, CoT may constrain the model to a fixed alignment between speech and text, which may limit the potential of E2E models of leveraging paralinguistic information. While this work does not explore such signals here, it highlights Direct prompting as a viable strategy and points toward future scenarios where richer speech annotations could further enhance E2E S2TT.
2 Methodology
We systematically compare how CoT and Direct prompting strategies scale in S2TT. To do so, we generate pseudo-labeled S2TT data (S2TTpl) and evaluate a series of models trained with varying amounts of it.
2.1 Translation pseudo-labeling
We generate pseudo-labeled data by translating the transcriptions of an ASR dataset (Figure 1). We use the same translation LLM employed as backbone for the S2TT model. To filter out generation errors or low-quality translations, we first score the samples with a Quality Estimation (QE) system. We use BLASER 2.0 [10] for this task because it provides reference-free, sentence-level translation scores that correlate strongly with human judgments. We compute the scores by evaluating the similarity of the transcription and translation with blaser_2_0_qe [^1] https://huggingface.co/facebook/blaser-2.0-qe and remove samples below $3.75$, which was chosen after inspecting samples in multiple languages. Despite estimating the quality of the translation, BLASER 2.0 is based on SONAR [11] embeddings, which have a shared representation space for speech and text across languages. For this reason, we next use a Language Identification (LID) system to discard possible targets in the wrong language. We use GlotLID v3 [12] to compute the probabilities of each of the $L$ considered languages $\bm{p}=(p_{1},...,p_{L})$, and remove samples in which the expected target language $p_{e}<0.5$.
2.2 Speech Translation Model
We build an LLM-based E2E S2TT system for our experimentation, following a similar approach as [13, 14]. Each speech utterance xx is first encoded with a self-supervised model $f_{\text{enc}}$, extracting continuous representations. These features are quantized into a sequence of discrete tokens: $\bm{s}=(s_{1},\ldots,s_{T})=q\bigl(f_{\text{enc}}(x)\bigr),\quad s_{t}\in V_{s}$, where $q(\cdot)$ is the quantization function and $V_{s}$ is the speech token vocabulary. To adapt the pretrained LLM for speech inputs, its original vocabulary $V_{o}$ is expanded to $V=V_{o}\cup V_{s}$, alongside its corresponding token embedding matrix $E=\bigl[E_{o};\,E_{s}\bigr]\in\mathbb{R}^{(|V|)\times d}$, where $E_{s}$ is randomly initialized.
Training is performed in two stages. In stage 1, the backbone LLM is frozen and only the token embedding layer is trained on speech utterances from the ASR corpora, using next-token prediction. In stage 2, the entire LLM is trained on the data described in § 3.1.
For the backbone model we use salamandraTA-7B-Instruct, [^2] https://huggingface.co/BSC-LT/salamandraTA-7b-instruct a highly multilingual LLM fine-tuned for translation and proficient in 35 European languages. The speech encoder is mHuBERT [^3] https://huggingface.co/slprl/mhubert-base-25hz from TWIST [15], together with its trained k-means quantizer. This encoder was chosen because: (1) it supports multilingual speech and (2) it downsamples to 25 Hz, halving the temporal resolution compared to standard HuBERTs [16]. Discrete speech tokens are obtained by clustering the 11th–layer representations into 500 groups.
CoT Prompt Template {audio} Transcribe in {src language} {transcription}
Translate to {tgt language}
{translation}
Direct Prompt Template {audio} Translate to {tgt language} {translation}
Fig. 2: Prompt templates (model outputs are in bold)
To fully leverage ASR and T2TT performance in S2TT, CoT models use prompts derived from the template shown in Figure 2. The same template is applied during inference, and generation is performed with beam search using five beams.
3 Experimental Setup
3.1 Data
Our training framework includes four types of datasets: ASR, T2TT, S2TT and pseudo-labeled S2TT (S2TTpl) in six languages: Catalan (ca), German (de), English (en), Spanish (es), French (fr), and Italian (it). Table 1 reports the amount of data per language. For ASR, we use the train splits of Common Voice 21.0 [17] and Multilingual LibriSpeech [18], totaling approximately 6,000 and 48,900 hours, respectively. For T2TT, we use Wikimedia [19] samples with 5 to 100 words, filtered using the same pipeline applied to pseudo-labeled S2TT data. For S2TT, we consider the train splits of Europarl-ST v1.1 [20] and CoVoST 2 [21], totaling around 630 and 1,600 hours, respectively. Finally, S2TTpl data is generated by translating all Common Voice 21.0 samples into all possible target languages. We evaluate our models in CoVoST 2 and Fleurs [22] test sets.
Table 1: Amount of data per language and total target tokens (considering transcriptions and translations as target in S2TT). The number of hours for S2TTpl corresponds to repeated utterances (one for each of the five target languages).
ca de en es fr it tgt tok.
Languages as Source (hours)
ASR 1,775 2,900 46,422 1,412 1,895 493 610M
S2TT 136 265 1,163 173 355 141 48M
S2TTpl 8,875 4,670 8,810 2,470 4,090 1,230 384M
Languages as Target (samples)
T2TT 289k 48k 76k 668k 508k 160k 134M
S2TT 289k 341k 585k 54k 50k 48k 48M
S2TTpl 1.35M 2.05M 2.63M 2.18M 2.10M 2.36M 384M
3.2 Experiments
We train two baseline models using all the ASR, T2TT and S2TT data described in § 3.1, with direct (DirectBase) and CoT (CoTBase) prompting.
To analyze the impact of scaling S2TT data, we incrementally add S2TTpl samples to each baseline, training 5 new systems with 20%, 40%, 60%, 80% and 100% of the full dataset respectively. We refer to these models as DirectaugXX and CoTaugXX, where xx is the percentage of pseudo-labeled data used. These models are trained independently, they are not checkpoints of a single experiment.
In the CoT setting, each speech–transcription pair in the pseudo-labeled data is repeated six times, one for each target language. We hypothesize this repetition may bias the model and harm the ASR performance, which is particularly critical for CoT prompting. To validate this, we train two variants of CoTaugXX: in CoTaugXX, the transcription step contributes to the loss, whereas in CoTaugXX it does not.
We evaluate the models using Word Error Rate (WER) for ASR, and with BLEU [23], computed with SacreBLEU [24],444signature: nrefs:1|case:mixed|eff:no|tok:13a|smooth: exp|version:2.5.1 as well as xcomet [25] [^5] https://huggingface.co/Unbabel/XCOMET-XL for T2TT and S2TT.
(a) S2TT
(b) ASR
(c) T2TT
Fig. 3: Performance on Fleurs (xcomet) when scaling the amount of S2TTpl data. Results are averaged across all the considered language directions.
3.3 Training Details
Both stage 1 and stage 2 last one epoch and employ AdamW optimizer, a cosine learning rate scheduler, and gradient clipping with a maximum norm of $1.0$.
In stage 1, we use a learning rate of $7\cdot 10^{-5}$, a warm up during the first $3\%$ of updates, and set a maximum sequence length of 1024. Tokens from different sequences are packed into full-length sequences to improve efficiency, following [13]. Training is run on 16 GPUs with a per-device batch size of 64, yielding an effective batch size of 256.
In stage 2, we increase the maximum sequence length to 2048 (without sequence packing) and lower the learning rate to $4\cdot 10^{-5}$, with warmup over the first $10\%$ of updates. We train on 16 GPUs with a per-device batch size of 32, resulting in an effective batch size of 512.
All experiments are implemented in the transformers library with DeepSpeed for distributed training, and run on NVIDIA H100 GPUs. To optimize memory usage, we apply mixed precision (bfloat16), gradient checkpointing, and Liger Kernel, which further improves efficiency and throughput.
4 Results
Initial CoTBase superiority  We first verify that our baselines reproduce trends reported in prior work [4]. As expected, Table 2 shows that the CoTBase outperforms the DirectBase across all languages, confirming the benefit of decomposing the task into transcription and translation in this scenario. The average baseline gap is approximately 5 BLEU points and 7 xcomet points.
Table 2: Baselines evaluation on FLEURS.
BLEU xcomet
[…]
DirectBase 21.04 22.80 30.32 80.6 79.7 86.0
CoTBase 26.39 29.76 33.24 88.0 87.2 88.6
Direct scales better than CoT  Figure 3(a) shows that the two CoTaug variants peak at 20% of the S2TTpl data and then degrade as more data is added. In contrast, Directaug consistently improves, achieving a better score with Directaug100. Although the peaks of CoTaug and CoTaug are still higher, Directaug demonstrates a steadier scaling trend, suggesting that this gap could be closed with additional data. These results indicate a potential advantage of direct prompting when sufficient S2TTpl data is available.
CoT degrades ASR as data scales  Figure 3(b) shows that WER remains stable with Directaug across all S2TTpl data scales, whereas both CoTaug and CoTaug gradually increase it by up to +7.4+7.4% and +13+13% relative to their baseline. These trends correlate with the steady decrease of S2TT performance, reflecting their critical reliance on the ASR sub-task. This is also observed at 2020% of S2TTpl data, where CoTaug reaches its minimum WER alongside its highest xcomet score. Additionally, CoTBase shows a WER 22% lower than its counterpart strategy. This confirms the findings of [8], indicating that direct S2TT data improve ASR performance.
Figure 3(c) shows that Direct performs slightly worse on T2TT, likely due to the lower proportion of this task in the training recipe (in Directaug, S2TTpl samples do not contain the T2TT sub-task). Nevertheless, performance for all three methods remains stable around +2 xcomet points. These results make Direct appealing, showing that ASR and T2TT remain stable as pseudo-labeled S2TT data scales.
[…]
Fig. 4: S2TT results evaluated on Fleurs test set with xcomet. The x-axis represents the percentage of the S2TTpl data with which the different Direct and CoT models were trained.
CoTaug lags behind CoTaug  Figures 3(a) and 3(b) support the hypothesis from § 3.2 that training the transcription step may negatively impact the model, as CoTaug consistently underperforms CoTaug in ASR and S2TT. Nonetheless, CoTaug also leads to some degradation. A possible explanation is that, by not training the transcription step, all S2TTpl data is treated as T2TT with speech as context, effectively reducing the proportion of the ASR task.
Results are consistent across languages  To assess whether these results hold across languages, Figure 4 shows S2TT performance for en, ca, and it. We selected these languages because en benefits from the the majority of ASR data and typically achieves strong translation performance due to the abundance of English-centric datasets. The other two are extremes in S2TTpl data quantity (see Table 1): ca has the most and it the least.
[…]
5 Conclusions and Future Directions
In this work, we systematically compare CoT and Direct prompting strategies for S2TT using pseudo-labeled data, which has become the standard approach for scaling this task. Our experiments show that CoT struggles to improve with more data, regardless of whether the transcription step is trained or not. In contrast, Direct prompting demonstrates a consistent scaling trend: although its absolute performance has not yet surpassed CoT, it steadily improves as more pseudo-labeled data is added, indicating strong potential for larger scale trainings.
These findings position Direct prompting as a promising strategy for future research. Beyond its favorable scaling behavior, Direct also offers practical advantages: it requires roughly half the computational resources of CoT and is simpler to implement. Moreover, the observed trends raise important questions about its limits: it remains to be validated whether Direct could not only match but eventually surpass CoT with an even larger scale of data, clarifying its potential as a primary strategy for S2TT training.
Finally, beyond the current setup, a particularly relevant direction is paralinguistic-aware translation. E2E S2TT systems are not constrained by the transcription bottleneck, which gives them the potential to produce richer translations that reflect linguistic cues uniquely present in speech, such as prosody. However, because current training largely relies on ASR and T2TT datasets, especially in CoT prompting, this capacity is unlikely to emerge naturally. Progress may therefore require training with direct translations annotated for speech-specific cues, and it would be valuable to investigate whether Direct prompting can better exploit the full richness of the speech signal compared to CoT.
6 Acknowledgements
This work is funded by the Ministerio para la Transformación Digital y de la Función Pública and Plan de Recuperación, Transformación y Resiliencia – Funded by EU – NextGenerationEU within the framework of the project Modelos del Lenguaje. FC and CEB acknowledge their AI4S fellowship within the “Generación D” initiative by Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent attraction (C005/24-ED CV1), funded by NextGenerationEU through PRTR.

### v2

@@
 The dominant trend for creating S2TT resources is to translate existing ASR corpora into target languages [17, 22].
-Following this approach, we construct a pseudo-labeled dataset through a translation pipeline (Figure 1), and train models with both Direct and CoT while scaling the amount of data.
+Following this approach, we construct a pseudo-labeled dataset using a translation pipeline (Figure 1), and train models with both Direct and CoT while scaling the amount of data.
 Our results demonstrate that Direct exhibits more stable scaling than CoT.
@@
 We generate pseudo-labeled data by translating the transcriptions of an ASR dataset (Figure 1).
-We use the same translation LLM employed as backbone for the S2TT model.
+We use the same translation LLM employed as backbone for the S2TT model (see §2.2).
 To filter out generation errors or low-quality translations, we first score the samples with a Quality Estimation (QE) system.
 We use BLASER 2.0 [4] for this task because it provides reference-free, sentence-level translation scores that correlate strongly with human judgments.
-We compute the scores by evaluating the similarity of the transcription and translation with blaser_2_0_qe [^1] https://huggingface.co/facebook/blaser-2.0-qe and remove samples below $3.75$, which was chosen after inspecting samples in multiple languages.
-Despite estimating the quality of the translation, BLASER 2.0 is based on SONAR [11] embeddings, which have a shared representation space for speech and text across languages.
-For this reason, we next use a Language Identification (LID) system to discard possible targets in the wrong language.
-We use GlotLID v3 [12] to compute the probabilities of each of the $L$ considered languages $\bm{p}=(p_{1},...,p_{L})$, and remove samples in which the expected target language $p_{e}<0.5$.
+We compute the scores by evaluating the similarity of the transcription and translation with blaser_2_0_qe [^1] https://huggingface.co/facebook/blaser-2.0-qe and remove samples below $3.75$, which was determined through inspection of samples across multiple languages.
+Although BLASER 2.0 estimates translation quality, it is based on SONAR [7] embeddings, which encode speech and text from multiple languages into a shared space.
+Thus, it does not explicitly enforce the intended target language.
+We therefore apply a Language Identification (LID) system, using GlotLID v3 [15], to compute language probabilities $\bm{p}=(p_{1},...,p_{L})$ and discard samples with expected language probability $p_{e}<0.5$.
 2.2 Speech Translation Model
@@
 Fig. 2: Prompt templates (model outputs are in bold)
-To fully leverage ASR and T2TT performance in S2TT, CoT models use prompts derived from the template shown in Figure 2.
+We use the prompts defined in Figure 2 for S2TT.
+We derive the templates from CoT for ASR and T2TT, which allows better integration of model capabilities from these tasks into S2TT.
 The same template is applied during inference, and generation is performed with beam search using five beams.
@@
 For T2TT, we use Wikimedia [21] samples with 5 to 100 words, filtered using the same pipeline applied to pseudo-labeled S2TT data.
-For S2TT, we consider the train splits of Europarl-ST v1.1 [20] and CoVoST 2 [21], totaling around 630 and 1,600 hours, respectively.
+For S2TT, we consider the train splits of Europarl-ST v1.1 [14] and CoVoST 2 [23], totaling about 630 and 1,600 hours, respectively.
 Finally, S2TTpl data is generated by translating all Common Voice 21.0 samples into all possible target languages.
@@
 Table 1: Amount of data per language and total target tokens (considering transcriptions and translations as target in S2TT).
-The number of hours for S2TTpl corresponds to repeated utterances (one for each of the five target languages).
+The number of S2TTpl hours reflects repeated utterances, as each utterance is paired with five target languages.
 ca de en es fr it tgt tok.
-Languages as Source (hours)
+Hours per source language
 ASR 1,775 2,900 46,422 1,412 1,895 493 610M
@@
 S2TTpl 8,875 4,670 8,810 2,470 4,090 1,230 384M
-Languages as Target (samples)
+Samples per target language
 T2TT 289k 48k 76k 668k 508k 160k 134M
@@
 We train two baseline models using all the ASR, T2TT and S2TT data described in §3.1, with direct (DirectBase) and CoT (CoTBase) prompting.
-To analyze the impact of scaling S2TT data, we incrementally add S2TTpl samples to each baseline, training 5 new systems with 20%, 40%, 60%, 80% and 100% of the full dataset respectively.
+To analyze the impact of scaling S2TT data, we incrementally add S2TTpl samples to each baseline, training 5 new systems with 20%, 40%, 60%, 80% and 100% of the pseudo-labeled dataset.
 We refer to these models as DirectaugXX and CoTaugXX, where xx is the percentage of pseudo-labeled data used.
@@
 In the CoT setting, each speech–transcription pair in the pseudo-labeled data is repeated six times, one for each target language.
-We hypothesize this repetition may bias the model and harm the ASR performance, which is particularly critical for CoT prompting.
+We hypothesize this repetition may harm the ASR performance, which is particularly critical for CoT prompting.
 To validate this, we train two variants of CoTaugXX: in CoTaugXX, the transcription step contributes to the loss, whereas in CoTaugXX it does not.
@@
 (c) T2TT
-Fig. 3: Performance on Fleurs (xcomet) when scaling the amount of S2TTpl data. Results are averaged across all the considered language directions.
+Fig.
+3: Performance on the Fleurs benchmark when scaling the amount of S2TTpl data.
+S2TT and T2TT are evaluated with xcomet ($\uparrow$) and ASR with WER ($\downarrow$).
+Note that $0\%$ corresponds to base variants and the rest to augXX.
+Results are averaged across all the considered language directions.
 3.3 Training Details
-Both stage 1 and stage 2 last one epoch and employ AdamW optimizer, a cosine learning rate scheduler, and gradient clipping with a maximum norm of $1.0$.
+Both stage 1 and stage 2 (see §2.2) last one epoch and employ AdamW optimizer, a cosine learning rate scheduler, and gradient clipping with a maximum norm of $1.0$.
 In stage 1, we use a learning rate of $7\cdot 10^{-5}$, a warm up during the first $3\%$ of updates, and set a maximum sequence length of 1024.
 Tokens from different sequences are packed into full-length sequences to improve efficiency, following [25].
-Training is run on 16 GPUs with a per-device batch size of 64, yielding an effective batch size of 256.
+Training is run on 16 GPUs with a per-device batch size of 16, yielding an effective global batch size of 256.
 In stage 2, we increase the maximum sequence length to 2048 (without sequence packing) and lower the learning rate to $4\cdot 10^{-5}$, with warmup over the first $10\%$ of updates.
@@
 This is also observed at 2020% of S2TTpl data, where CoTaug reaches its minimum WER alongside its highest xcomet score.
-Additionally, CoTBase shows a WER 22% lower than its counterpart strategy.
+Additionally, CoTBase shows a 2-point WER reduction compared to its counterpart strategy.
 This confirms the findings of [22], indicating that direct S2TT data improve ASR performance.
@@
 […]
-Fig. 4: S2TT results evaluated on Fleurs test set with xcomet. The x-axis represents the percentage of the S2TTpl data with which the different Direct and CoT models were trained.
+Fig.
+4: S2TT results evaluated on Fleurs test set with xcomet.
+The x-axis represents the percentage of the S2TTpl data with which the different Direct and CoT models were trained.
+Note that $0\%$ corresponds to base variants and the rest to augXX.
 CoTaug lags behind CoTaug  Figures 3(a) and 3(b) support the hypothesis from §3.2 that training the transcription step may negatively impact the model, as CoTaug consistently underperforms CoTaug in ASR and S2TT.